# Coherent Audio-Visual Editing via Conditional Audio Generation Following Video Edits

### v1

@@ 4.3 Main results
Figure 6 presents the performance of all editing methods. In both plots, the horizontal axis represents the degree of structural preservation (LPAPS). In the left plot, the vertical axis shows the audio-visual alignment between the target audio and video (IB-AV), while, in the right plot, it shows the text fidelity of the target audio (IB-TA). Figure 7 shows examples of the generated audio and video. For more examples, please refer to the supplementary materials.
Here are several observations we obtained through this experiment:
  • Benefit of sequential approach. The sequential editing methods achieved significantly better performance in audio-visual alignment as they refer to the edited video when generating the edited audio. In contrast, the independent editing methods achieved relatively good performance in terms of text fidelity, but resulted in poor audio-visual alignment due to the lack of a mechanism for aligning edited audio and video.
@@ 4.4 Ablation study
Figure 8: Ablation study.
We also performed an ablation study to clarify how much each newly introduced component contributes to enhancing the quality of the edited audio. Specifically, we examine our method to remove the SynchFormer feature modulation (SM) described in Section 3.3, the temporal masking (TM) in Section 3.4, and the adaptive conditioning (AC) in Section 3.5. In the case without AC, we used a fixed level of detail across all samples and varied it to control the overall degree of structure preservation.

### v2

@@ 4.3 Main results
Figure 6 summarizes the performance of all editing methods. In both plots, the horizontal axis indicates structural preservation (LPAPS). The left plot reports audio-visual alignment between the target audio and video (IB-AV), and the right plot reports target-audio text fidelity (IB-TA). Figure 7 shows qualitative examples. More examples are provided in the supplementary materials.
Here are several observations we obtained through this experiment:
  • Benefit of sequential approach. The sequential editing methods achieved significantly better performance in audio-visual alignment as they leverage the latest video-to-audio architecture when generating the edited audio. In contrast, the independent editing methods achieved resulted in poor audio-visual alignment due to the lack of a mechanism for aligning edited audio and video. AvED also failed to achieve high audio-visual alignment due to the difficulty of jointly editing both audio and video.
4.4 Subjective test
We also conducted a subjective test to evaluate the perceptual quality of the edited videos. In this test, each rater first watches and listens to the source video and then views each edited version. The rater is asked to rate the quality of each edited video on a scale from 1 (poor) to 5 (excellent) in terms of the following three aspects: audio-text fidelity, audio-visual alignment, and audio structure preservation. In total, eleven raters participated in the test, and each rater evaluated six different videos.
Table 1 shows the results of the subjective test. Our method achieved the highest scores across all metrics. Consistent with the main evaluation results (Fig. 6), the proposed method and V2A achieved high audio-text fidelity and audio-visual alignment. In contrast, for audio structure preservation, unlike the main evaluation results, ZETA and AvED showed low scores. This suggests that, when the edited audio is experienced together with the video, perceived consistency with the video strongly influences subjective ratings.
audio-text fidelity audio-visual alignment audio structure perservation
VACE $\rightarrow$ ZETA 3.0 ($\pm$ 0.5) 2.6 ($\pm$ 0.7) 2.5 ($\pm$ 0.7)
VACE $\rightarrow$ V2A 3.2 ($\pm$ 0.6) 3.3 ($\pm$ 0.6) 3.2 ($\pm$ 0.5)
AvED 2.7 ($\pm$ 0.4) 2.5 ($\pm$ 0.6) 3.1 ($\pm$ 0.4)
VACE $\rightarrow$ Ours 3.7 ($\pm$ 0.5) 3.8 ($\pm$ 0.3) 3.9 ($\pm$ 0.2)
Table 1: Subjective evaluation results (mean score $\pm$ 95% confidence interval). Bold indicates the best score for each metric.
@@ 4.5 Ablation study
Figure 8: Ablation study. AC, SM, and TM denote adaptive conditioning, Synchformer feature modulation, and temporal masking, respectively.
We also performed an ablation study to clarify how much each newly introduced component contributes to enhancing the quality of edited audio. Specifically, we examine our method to remove the Synchformer feature modulation (SM) described in Section 3.3, the temporal masking (TM) in Section 3.4, and the adaptive conditioning (AC) in Section 3.5. In the case without AC, we used a fixed level of detail in all samples and varied it to control the overall degree of structure preservation.